# Lost in Transcription, Found in Distribution Shift: Demystifying Hallucination in Speech Foundation Models

### v1

@@ Abstract
Speech foundation models trained at a massive scale, both in terms of model and data size, result in robust systems capable of performing multiple speech tasks, including automatic speech recognition (ASR). These models transcend language and domain barriers, yet effectively measuring their performance remains a challenge. Traditional metrics like word error rate (WER) and character error rate (CER) are commonly used to evaluate ASR performance but often fail to reflect transcription quality in critical contexts, particularly when detecting fabricated outputs. This phenomenon, known as hallucination, is especially concerning in high-stakes domains such as healthcare, legal, and aviation, where errors can have severe consequences. In our work, we address this gap by investigating hallucination in ASR models. We examine how factors such as distribution shifts, model size, and model architecture influence the hallucination error rate (HER), a metric we introduce to quantify hallucinations. Our analysis of 20 ASR models reveals three key insights: (1) High WERs can mask low hallucination rates, while low WERs may conceal dangerous hallucinations. (2) Synthetic noise, both adversarial and common perturbations like white noise, pitch shift, and time stretching, increase HER. (3) Distribution shift correlates strongly with HER ($\alpha=0.91$). Our findings highlight the importance of incorporating HER alongside traditional metrics like WER to better assess ASR model performance, particularly in high-stakes domains.
@@ 1 Introduction
Automatic Speech Recognition (ASR) systems have become fundamental to various applications, including personal assistants, automated customer service, and transcription tools used in fields such as education, healthcare, and law Zhang et al. (2023); Adedeji et al. (2024). These systems have seen remarkable improvements in recent years Arriaga et al. (2024); Radford et al. (2022); Communication et al. (2023), with state-of-the-art models demonstrating their capabilities across diverse datasets and languages Shakhadri et al. (2025). However, the evaluation of ASR performance remains largely dependent on word and character error rate (WER/CER). The primary limitation of WER and CER is their dependence on token-level overlapping, which focuses on matching individual words or characters without considering the overall semantic aspect of the transcription. This could result in misleading evaluations, as a high WER/CER does not necessarily indicate poor outputs in all cases. In addition, these metrics fall short in capturing more subtle semantic failures which aren’t typically caught without human verification, such as hallucinations.
@@
The existing literature on hallucination detection in ASR models is confined to a single model Frieske and Shi (2024); Serai et al. (2021); Koenecke et al. (2024a); Barański et al. (2025); Ji et al. (2023) or test setting Kim et al. (2024), highlighting a significant research gap for a systematic investigation across different supervision paradigms, test domains, and conditions.
@@
Outline. In Section 2, we review prior work related to ASR evaluation and hallucination detection. Section 3 outlines our proposed methodology. Section 4 presents our experimental setup. Finally, Section 5 discusses the results and implications of our findings, outlining directions for future research. We conclude our work in 6 and provide limitations in 7.
@@
whisper-large-v3 ripped ocean jumper. Thank you. 100 F T T Home Environment
hf-seamless-m4t-large probably i i had asthma What about Erasmus? 100 F T T Medical
hf-seamless-m4t-large it can’t be done I’m going to start with the first one. 180 F T T Legal
hf-seamless-m4t-medium News of this album coincides with John Robert’s departure from the band. The first is the oldest, which is also the oldest, the oldest, the oldest, the oldest, the oldest. 130 F T T Accented Speech
@@ 2 Related Work
Recent studies have extensively investigated hallucination in text generated by large language models (LLMs), identifying it as a prevalent phenomenon Huang et al. (2025); Bai et al. (2024); Yao et al. (2023); Jiang et al. (2024); Maynez et al. (2020); Parikh et al. (2020); Ji et al. (2023); Mittal et al. (2024); Filippova (2020). This issue has also been observed in audio foundation models Sahoo et al. (2024). Furthermore, research suggests that pretraining language models for predictive accuracy inherently predispose them to hallucination, even under ideal conditions where the training data is entirely factual Kalai and Vempala (2024).
However, few studies explore hallucination evaluation and detection in automatic speech recognition (ASR) systems, with most research focusing on Whisper, a semi-supervised model. For instance, Koenecke et al. (2024b) analyze the Aphasia dataset and report that while Whisper’s overall hallucination rate is 1%, 40% of these hallucinations contain violent or harmful content. Similarly, Kim et al. (2024) demonstrate that Whisper hallucinates at significantly higher rates under low signal-to-noise ratio (SNR) conditions, and observe a 20% increase in hallucinations at -4 dB and -2 dB SNRs. Prior work by Serai et al. (2021) propose augmenting models with hallucinated transcripts to improve performance, while Frieske and Shi (2024) introduce a perturbation-based evaluation method using automatic metrics such as word error rate (WER), perplexity, and cosine similarity. Barański et al. (2025) develop a filtered Bag of Hallucinations (BoH) for detection, and reveal that hallucinations in Whisper correlate strongly with training data biases (e.g., phrases like “Thank you for watching” linked to YouTube content).
@@ 4.1 Datasets
To simulate challenging acoustic conditions and evaluate WER and HER under adversarial scenarios, we apply various types of synthetic perturbations to speech inputs. These include an adversarial dataset featuring modified utterances with adversarial noise at varying radii (0.04 and 0.015) and Room Impulse Response (RIR) noise, primarily aimed at adversarially attacking the ASR models Olivier and Raj (2022). Additionally, we evaluate model robustness under challenging conditions by applying a range of general audio perturbations—including noise addition, time stretching, and pitch shift—to 1,000 randomly sampled audio clips from a mixture of domain-specific datasets. These perturbations are commonly used as augmentation techniques to simulate real-world variability and stress-test the models. We compare the performance on perturbed speech with that of non-perturbed speech to quantify the impact of these distortions. Full details about the perturbation methods, including parameters and implementation, are provided in Appendix A.1.2 (Table 5). Additionally, comprehensive information about the datasets used in this study can be found in Appendix A.1.1 (Table 4).
@@ 4.2 Models
  • Decoder-only models: Qwen2Audio Chu et al. (2024) and SpeechLLM Rajaa and Tushar , optimized for text generation and audio-language alignment.
  • Encoder-decoder models: Whisper Radford et al. (2022) (10 variants), DistilWhisper Gandhi et al. (2023) (4 variants), and SeamlessM4T Communication et al. (2023) (2 variants), designed for multilingual transcription, translation, and speech-to-text tasks.
The selected models vary in size (39M to 7B parameters), depth (4 to 32 layers), and training paradigms (supervised, self-supervised, semi-supervised). Full specifications, including architectural details, training data, and hyperparameters, are provided in Appendix A.1.3 (Table 6).
@@ 5 Results
5.1 Hallucination Error Detection
We assess the ASR outputs of large language models (LLMs) by classifying them into various error categories. Specifically, we use GPT-4o mini to identify the types of errors in ASR outputs. The evaluation process includes both coarse-grained and fine-grained error classifications, as detailed in Section 3.2. The prompts used for both evaluations are shown in Figures 5 and 6.
Coarse-grained and Fine-grained Evaluation.
Figure 8 in the appendix illustrates the error distributions across coarse-grained and fine-grained hallucination categories. Our results demonstrate strong alignment between both levels, indicating consistent classification of hallucinations. Among non-hallucination errors, phonetic errors dominate across most datasets (see Appendix Table 9). However, in Primock57, language errors prevail, likely due to its specialized medical terminology. This aligns with Ferrando et al. (2024), who emphasize language models’ struggles with domain-specific named entities. This is also reflected in the third example provided in Table 1.
Agreement with Human Evaluation and Heuristic Baseline.
To validate our approach, we compare model-to-human and human-to-human agreement scores using a coarse-grained prompt. Our results demonstrate strong human-to-human raw agreement (0.71), indicating consistency. Additionally, we observe good agreement (0.6) between human annotations and GPT-4o-mini’s coarse-grained output, suggesting that the model aligns reasonably well with human judgments. We further evaluate the agreement between human and model classifications against a heuristic baseline proposed by Frieske and Shi (2024). Their method is based on a cosine similarity threshold of 0.2, alongside a WER threshold of 30 and a Flan-T5 Chung et al. (2024) perplexity threshold of 200. However, as shown in Table 3, this heuristic achieves significantly lower agreement scores: 0.1 with GPT-4o mini and 0.14 with Gemini-2.0-flash-001. These results highlight the limitations of purely heuristic-based approaches compared to our method, which better captures the more fine-grained aspects like hallucination.
Evaluation Pair Agreement Score ($\pm$)
Human & Human 2 0.71 $\pm$ 0.01
Human & heuristic 0.00 $\pm$ 0.00
Human 2 & GPT_coarsegrained 0.60 $\pm$ 0.01
Human 2 & Gemini-2.0 0.59 $\pm$ 0.01
GPT_coarsegrained & heuristic 0.10 $\pm$ 0.00
GPT_coarsegrained & Gemini-2.0 0.78 $\pm$ 0.01
heuristic & Gemini-2.0 0.14 $\pm$ 0.01
Table 3: Raw agreement scores between different hallucination evaluation methods.
@@ 5.2 Errors Under Distributional Shifts
Natural Shift.
@@
Synthetic Shift.
Under synthetic shift, we experiment with two configurations: (a) adversarial perturbations and (b) common perturbations. Our experiments reveal that adversarial attacks cause the most significant degradation in HER, with adversarial datasets showing the highest HER values across all models, exceeding the degradation observed under natural shift baselines (see Table 2). In contrast, random perturbations (such as white noise, pitch shifts, and time stretching) result in more moderate impacts. Notably, self- and semi-supervised models like whisper and seamless demonstrate consistent vulnerability to structured perturbations. For instance, pitch shifts and time stretching lead to a substantial increase of approximately (242%) in both WER and HER across these models, while white noise causes smaller degradations of approximately (142%). Interestingly, the supervised wav2vec2 model exhibits non-uniform behavior, where the impact on WER and HER is similar across all perturbations. Furthermore, it is noteworthy that HER increased by 50% in the wav2vec2 model, which is considerably less than the sharp increase observed in whisper and seamless models, highlighting a notable contrast in the robustness of these models to random synthetic shifts against more targeted shifts.
@@
5.3 Impact of Model Architecture and Scale on Error Rates
Model Type.
We hypothesize that model architecture plays a critical role in influencing error rates, with these effects further modulated by the scale and diversity of the training data. As shown in Appendix Table 8, our findings indicate that encoder-only models Wav2vec2-large-xlsr-53-english and Hubert-large-ls960-ft exhibit the lowest HER/WER ratios across multiple datasets, particularly when compared to models like Qwen2-Audio-7B and hf-seamless-m4t-large. This suggests that encoder-only models are less prone to hallucinations relative to their overall errors, likely due to robust training or architectural advantages.
Model Size.

### v2

@@ Abstract
Speech foundation models trained at a massive scale, both in terms of model and data size, result in robust systems capable of performing multiple speech tasks, including automatic speech recognition (ASR). These models transcend language and domain barriers, yet effectively measuring their performance remains a challenge. Traditional metrics like word error rate (WER) and character error rate (CER) are commonly used to evaluate ASR performance but often fail to reflect transcription quality in critical contexts, particularly when detecting fabricated outputs. This phenomenon, known as hallucination, is especially concerning in high-stakes domains such as healthcare, legal, and aviation, where errors can have severe consequences. In our work, we address this gap by investigating hallucination in ASR models. We examine how factors such as distribution shifts, model size, and model architecture influence the hallucination error rate (HER), a metric we introduce to quantify hallucinations. Our analysis of over 20 ASR models reveals three key insights: (1) High WERs can mask low hallucination rates, while low WERs may conceal dangerous hallucinations. (2) Synthetic noise, both adversarial and common perturbations like white noise, pitch shift, and time stretching, increase HER. (3) Distribution shift correlates strongly with HER ($\alpha=0.91$). Our findings highlight the importance of incorporating HER alongside traditional metrics like WER to better assess ASR model performance, particularly in high-stakes domains.
@@ 1 Introduction
Automatic Speech Recognition (ASR) systems have become fundamental to various applications, including personal assistants, automated customer service, and transcription tools used in fields such as education, healthcare, and law Zhang et al. (2023); Adedeji et al. (2024). These systems have seen remarkable improvements in recent years Arriaga et al. (2024); Radford et al. (2022); Communication et al. (2023), with state-of-the-art models demonstrating their capabilities across diverse datasets and languages Shakhadri et al. (2025). However, the evaluation of ASR performance remains largely dependent on word and character error rate (WER/CER). The primary limitation of WER and CER is their dependence on token-level overlapping, which focuses on matching individual words or characters without considering the overall semantic aspect of the transcription. This could result in misleading evaluations, as a high WER/CER does not necessarily indicate poor outputs in all cases. In addition, these metrics fall short in capturing more subtle semantic failures which are not typically caught without human verification, such as hallucinations.
@@
The existing literature on hallucination detection in ASR models is confined to a single model Frieske and Shi (2024); Serai et al. (2021); Koenecke et al. (2024a); Barański et al. (2025); Ji et al. (2023) or test setting Kim et al. (2024a), highlighting a significant research gap for a systematic investigation across different supervision paradigms, test domains, and conditions.
@@
Outline. In Section 2, we review prior work related to ASR evaluation and hallucination detection. Section 3 outlines our proposed methodology. Section 4 presents our experimental setup. Finally, Section 5 discusses the results and implications of our findings, outlining directions for future research. We conclude our work in Section 6 and provide limitations in Section 7.
@@
hf-seamless-m4t-large probably i i had asthma The song is Erasmus. 100 F T T Primock57 (medical)
hf-seamless-m4t-large it can’t be done I’m going to start with the first one. 200 F T T Superemecourt (Legal)
@@
hubert god i’m forty five god he whisper i’m forty five 40 F T T Primock57 (medical)
hubert yeah in a handsome coat 400 F T T AMI (meetings)
wav2vec today yeah she did yes 150 F T T Primock57 (medical)
wav2vec2 another dog secretary show i love her dog’s secretary show 100 F T T BERSt (Noisy)
@@ 2 Related Work
Recent studies have extensively investigated hallucination in text generated by large language models (LLMs), identifying it as a prevalent phenomenon Huang et al. (2025); Bai et al. (2024); Yao et al. (2023); Jiang et al. (2024); Maynez et al. (2020); Parikh et al. (2020); Ji et al. (2023); Mittal et al. (2024); Filippova (2020). This issue has also been observed in audio foundation models Sahoo et al. (2024). Furthermore, research suggests that pretraining language models for predictive accuracy inherently predisposes them to hallucination, even under ideal conditions where the training data is entirely factual Kalai and Vempala (2024).
However, few studies explore hallucination evaluation and detection in automatic speech recognition (ASR) systems, with most research focusing on Whisper, a semi-supervised model. For instance, Koenecke et al. (2024b) analyzes the Aphasia dataset and reports that while Whisper’s overall hallucination rate is 1%, 40% of these hallucinations contain violent or harmful content. Similarly, Kim et al. (2024a) demonstrates that Whisper hallucinates at significantly higher rates under low signal-to-noise ratio (SNR) conditions, and observes a 20% increase in hallucinations at -4 dB and -2 dB SNRs. Prior work by Serai et al. (2021) proposes augmenting models with hallucinated transcripts to improve performance, while Frieske and Shi (2024) introduces a perturbation-based evaluation method using automatic metrics such as word error rate (WER), perplexity, and cosine similarity. Barański et al. (2025) develops a filtered Bag of Hallucinations (BoH) for detection, and reveals that hallucinations in Whisper correlate strongly with training data biases (e.g., phrases like “Thank you for watching” linked to YouTube content).
@@ 4.1 Datasets
To simulate challenging acoustic conditions and evaluate WER and HER under adversarial scenarios, we apply various types of synthetic perturbations to speech inputs. These include an adversarial dataset featuring modified utterances with adversarial noise at varying radii (0.04 and 0.015) and Room Impulse Response (RIR) noise, primarily aimed at adversarially attacking the ASR models Olivier and Raj (2022). Additionally, we evaluate model robustness under challenging conditions by applying a range of general audio perturbations—including noise addition, time stretching, pitch shift, cross-lingual noise, distortion, echo, pub noise, and reverberation—to 1,000 randomly sampled audio clips from a mixture of domain-specific datasets. These perturbations are commonly used as augmentation techniques to simulate real-world variability and stress-test the models. We compare the performance on perturbed speech with that of non-perturbed speech to quantify the impact of these distortions. Comprehensive information about the datasets used in this study can be found in Appendix A.2.1 (Table 10).
@@ 4.2 Models
  • Encoder-decoder models: Whisper Radford et al. (2022) (10 variants), DistilWhisper Gandhi et al. (2023) (4 variants), and SeamlessM4T Communication et al. (2023) (2 variants) are designed for multilingual transcription, translation, and speech-to-text tasks. We also include Qwen2Audio Chu et al. (2024) and SpeechLLM Rajaa and Tushar , which are optimized for text generation and audio-language alignment. Additionally, we incorporate Canary Harper et al.  [^3] https://huggingface.co/nvidia/canary-1b, which uses FastConformer Rekesh et al. (2023) as its encoder and a transformer decoder.
In addition to these, we also include open-weight transducer-based models. Specifically, we add a model trained with transducer decoder loss (parakeet-tdt-1.1b) and a token-and-duration transducer Xu et al. (2023) (parakeet-tdt-1.1b), both from the NVIDIA-NeMo ASR repository Harper et al. . In summary, the selected models vary in size (39M to 7B parameters), depth (4 to 32 layers), and training paradigms (supervised, self-supervised, and semi-supervised). Full specifications, including architectural details, training data, and hyperparameters, are provided in Appendix A.2.2 (Table 11).
@@ 5 Results
5.1 Hallucination Error Detection
We assess the ASR outputs of large language models by classifying them into various error categories. Specifically, we use GPT-4o mini to identify the types of errors in ASR outputs. The evaluation process includes both coarse-grained and fine-grained error classifications, as detailed in Section 3.2. The prompts used for both evaluations are shown in Figures 5 and 6.
Verbalized Confidence Scores
Accurate error classification alone does not always provide insight into the model’s own uncertainty or reliability. To address this, we also assess the self-reported confidence of the LLM in its classifications. Inspired by Yang et al. (2024) and other similar works Tian et al. (2023); Xiong et al. (2024), we prompt GPT-4o-mini to provide explicit confidence scores (on a scale of 1–10) for each classification decision. This approach allows us to understand not just what the model predicts, but how confident it is in those predictions, offering an additional dimension for evaluating model performance and robustness. In this experiment, we take 500 examples, and after each classification, we simply ask the model in a multiturn conversation: “How confident are you about the classification on a scale of 1–10? Confidence:”. We report the resulting confidence scores in Tables 2 and 3 for fine-grained and coarse-grained evaluations, respectively.
The results show that GPT-4o-mini’s self-assessed confidence remains consistently high across most categories. In the fine-grained evaluation, the “Hallucination Error” (HE) and “Oscillation Error” (OE) classes achieve nearly perfect confidence scores, with median and 25%ile values of 10.0. The “No Error” (NE) predictions also exhibit strong confidence, with an overall average of 8.81 across fine-grained classes, highlighting robust certainty in identifying transcription errors.
Metric HE LE NE OE PE Overall
Mean 9.56 7.33 9.77 10.0 7.37 8.81
Median 10.0 7.0 10.0 10.0 7.0 8.80
1%ile 8.00 7.00 7.21 10.0 7.00 7.84
5%ile 8.0 7.0 8.1 10.0 7.0 8.02
10%ile 8.0 7.0 10.0 10.0 7.0 8.40
25%ile 9.0 7.0 10.0 10.0 7.0 8.60
Table 2: Self-verbalized confidence scores (1–10) of the LLM classifier for each fine-grained error class—Hallucination Error (HE), Language Error (LE), No Error (NE), Oscillation Error (OE), and Phonetic Error (PE)—along with overall average scores.
In contrast, the “Language Error” (LE) and “Phonetic Error” (PE) classes show slightly lower confidence, with mean values around 7.3 and narrower interquartile ranges, reflecting more nuanced or ambiguous cases.
Metric HE NE NHE Overall
Mean 9.77 9.20 7.09 8.69
Median 10.0 10.0 7.0 9.00
1%ile 8.0 7.0 1.0 5.33
5%ile 8.0 7.95 7.00 7.65
10%ile 9.0 8.0 7.0 8.00
25%ile 10.0 8.0 7.0 8.33
Table 3: Self-verbalized confidence scores (1–10) of the LLM classifier for each coarse-grained error class—Hallucination Error (HE), No Error (NE), and Non-Hallucination Error (NHE)—along with overall average scores.
The coarse-grained evaluation mirrors these trends, with “Hallucination Error” and “No Error” classes maintaining high confidence (mean $\geq$ 9.2), while the “Non-Hallucination Error” (NHE) class has a lower mean of 7.09 and a notably low 1%ile of 1.0, indicating occasional uncertainty in ambiguous scenarios. Despite this, the overall average for the coarse-grained setting is 8.69, confirming the reliability of these scores.
Overall, these findings demonstrate that GPT-4o-mini not only provides reliable classification outputs but also accurately calibrates its confidence based on the error type. The high overall averages further validate the robustness of the LLM-based evaluation framework in discerning and categorizing transcription errors.
Coarse-grained and Fine-grained Evaluation.
Figure 8 in the appendix illustrates the error distributions across coarse-grained and fine-grained hallucination categories. Our results demonstrate strong alignment between both levels, indicating consistent classification of hallucinations. Among non-hallucination errors, phonetic errors dominate across most datasets (see Appendix Table 15). However, in Primock57, language errors prevail, likely due to its specialized medical terminology. This aligns with Ferrando et al. (2024), who emphasize language models’ struggles with domain-specific named entities. This is also reflected in the second example provided in Table 1.
Agreement with Human Evaluation and Heuristic Baseline.
To validate our approach, we compare model-to-human and human-to-human agreement scores using a coarse-grained prompt. Our results demonstrate strong human-to-human raw agreement (0.71), indicating consistency. Additionally, we observe good agreement (0.6) between human annotations and GPT-4o-mini’s coarse-grained output, suggesting that the model aligns reasonably well with human judgments.
Evaluation Pair Agreement Score ($\uparrow$)
Human - Human 0.71 $\pm$ 0.01
Human - Heuristic 0.00 $\pm$ 0.00
Human - GPT 0.60 $\pm$ 0.01
Human - Gemini 0.59 $\pm$ 0.01
GPT - Heuristic 0.10 $\pm$ 0.00
GPT - Gemini 0.78 $\pm$ 0.01
Heuristic - Gemini 0.14 $\pm$ 0.01
Table 4: Agreement scores ($\uparrow$) between different hallucination evaluation methods. Here, GPT-4o-mini is referred to as GPT, and Gemini-2.0-flash-001 is referred to as Gemini. All evaluations are coarse-grained.
We further evaluate the agreement between human and model classifications against a heuristic baseline proposed by Frieske and Shi (2024). Their method is based on a cosine similarity threshold of 0.2, alongside an WER threshold of 30 and a Flan-T5 Chung et al. (2024) perplexity threshold of 200. However, as shown in Table 4, this heuristic achieves significantly lower agreement scores: 0.1 with GPT-4o mini and 0.14 with Gemini-2.0-flash-001. These results highlight the limitations of purely heuristic-based approaches compared to our method, which better captures the more fine-grained aspects like hallucination.
@@ 5.2 Errors Under Distributional Shifts
We analyze how ASR performance degrades under distributional shifts—when models are exposed to domain conditions different from their training data or domains where they outperform the human baseline (referred to as the source domain) [^6] Here, “domain” is loosely defined; for some models, the source or training data is unknown. In those cases, we treat the source domain as the data on which the error is nearly zero, such as LibriSpeech.. For natural (domain-specific) shifts, we quantify the distribution shift and measure the resulting degradation in both WER and hallucination errors (HER), with results shown in Figure 2 and detailed breakdowns in Appendix A.4. For synthetic shifts (e.g., adversarial and random perturbations), we directly evaluate model robustness. In what follows, we describe these two shift types—natural and synthetic—separately.
Natural Shift
@@
Synthetic Shift
Under synthetic shift, we experiment with two configurations: (a) adversarial perturbations and (b) common perturbations. Our experiments reveal that adversarial attacks cause the most significant degradation in HER, with adversarial datasets showing the highest HER values across all models, exceeding the degradation observed under natural shift baselines (see Table 5). In contrast, random perturbations (such as white noise, pitch shifts, and time stretching) result in more moderate impacts. Notably, self- and semi-supervised models like whisper and seamless demonstrate consistent vulnerability to structured perturbations. For instance, pitch shifts and time stretching lead to a substantial increase of approximately (242%) in both WER and HER across these models, while white noise causes smaller degradations of approximately (142%). Interestingly, the supervised wav2vec2 model exhibits non-uniform behavior, where the impact on WER and HER is similar across all perturbations. Furthermore, it is noteworthy that HER increased by 50% in the wav2vec2 model, which is considerably less than the sharp increase observed in whisper and seamless models, highlighting a notable contrast in the robustness of these models to random synthetic shifts against more targeted shifts. We provide results for additional perturbations in Appendix A.4.2 Table 8.
@@
5.3 Impact of Architecture and Scale
In this section, we analyze how model architecture and scale influence hallucination and transcription errors. We specifically examine how different model types—encoder-only, encoder-decoder, and transducer-based—affect hallucination error rates, as well as how scaling within a model family (such as whisper) impacts overall performance.
Model Type
We hypothesize that model architecture plays a critical role in influencing error rates, with these effects further modulated by the scale and diversity of the training data. As shown in Appendix Table 14, our findings indicate that encoder-only models Wav2vec2-large-xlsr-53-english and Hubert-large-ls960-ft exhibit the lowest HER/WER ratios across multiple datasets, particularly when compared to models like Qwen2-Audio-7B and hf-seamless-m4t-large. This suggests that encoder-only models are less prone to hallucinations relative to their overall errors.
Model Size